OpenStudy (anonymous):

solve 1: (-mn8)2 2: (3x3y)2(-4x2y4)3 3: 10 times x to the 4th power times y squared all divided by -5 x to the 3rd times y squared. The entire quantity is raised to the 3rd power.

OpenStudy (anonymous):

help

OpenStudy (anonymous):

11. is that squared?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok square everything in the ( )

OpenStudy (anonymous):

-m^2n^264 -64m^2n^2

OpenStudy (anonymous):

2. (3x3y)^2 will give (9xy)^2=9^2x^2y^2==81x^2y^2 (-4x^2y^4)^3 gives (-4^3)(x^2*3)(y^4*3)= -64x^6y^12 Now multiply (81*-64)(x^2+x^6)(y^2+y^12)= -5184x^8y^14

OpenStudy (anonymous):

3. 10x^4(y^2)/(-5x)^3y^2 the y^2 cancel because theres one in both the numerator and the denominator this leaves: 10x^4/-125x^3 reduce and you get-2x/25 10x^4/(-5^3)x^3=10x^4/-125x^4 simplified leaves -2x/25